Use the Heine-Borel Theorem to prove the Bolzano-Weierstrass The- orem.
Answer #1

ANSWER :

Proof :  

Proof of Bolzano - Weierstrass Theorem from Heine - Borel Theorem,

Let \left \{ a_{n} \right \} be a bounded sequences IR. Then a_{n} \epsilon \left [ c,d \right ] for some c,d \epsilon IR \, \, \, c,d To prove applying contradiction .

Assume \left \{ a_{n} \right \} contains no convergend subsequence ln \left [ c,d \right ] , then \forall x\, \epsilon \left [ c.d \right ],\exists\, \, \epsilon > 0 such that  B_{\epsilon }(x)=\left ( x-\epsilon ,x\rightarrow \epsilon \right )

Contains finitely many a_{n} .

Otherwise if there excused an x such that \left | x-a_{n} \right |< \epsilon\, \, \, \forall \rightarrow 0 and finitely many a_{n} .

\Rightarrowa_{n} would have a convergent subsequence contaning.

Then,

\Rightarrow \bigcup _{x\epsilon [c,d]}B_{\epsilon }(x)\supset [c,d]

Then by HBT a finite subcover such that ,

\bigcup_{i=1}^h{}B_{\epsilon }(xi)\supset [c,d]

Therefore, \left \{ a_{n} \right \} is contained subcover.

There we conclude that

\left \{ a_{n} \right \} has finitely marry.

Memvers \, \, bounded \Rightarrow \Leftarrow

Here, \left \{ a_{n} \right \}ln\, \, IR has a convergent subsequence ln R.